Integrate ∫dy/(y(y^2 + 1))


1
Expert's answer
2021-09-30T00:16:57-0400

Let us integrate "\int\frac{dy}{y(y^2 + 1)}:"

"\int\frac{dy}{y(y^2 + 1)}=\int(\frac{1}{y}-\frac{y}{y^2+1})dy
=\ln|y|-\frac{1}{2}\int(\frac{d(y^2+1)}{y^2+1})
\\=\ln|y|-\frac{1}{2}\ln(y^2+1)+C."